# Scaling Law with Learning Rate Annealing

### v1

@@ 4 Takeaways: Experimental Findings Verification and Explanation
Recently, it has been demonstrated that WSD LRS (Hu et al., 2024) and multi-step cosine LRS (DeepSeek-AI, 2024) yield a lower loss compared to the typical cosine LRS. We also prove this by the experiments (refer to the ground-truth loss in Fig. 3(b)3(c)3(d)). We validate and elucidate this finding using our proposed scaling law with LR annealing. In Fig. 10, we depict the learning rate (on the left) and the predicted loss drop (on the right) for different LRS using our derived equation. The figures suggest that for multi-step cosine or WSD LRS, the negative $S_{2}$-item ($-C\cdot S_{2}$) is slightly higher than that of cosine LRS. However, the $S_{1}$-item ($A\cdot S_{1}^{-\alpha}$) is significantly lower than that in cosine LRS. More simply to say, both WSD LRS and multi-step cosine LRS aim to employ a strategy that marginally reduces $S_{2}$ but substantially increases $S_{1}$, leading to an overall decrease in validation loss.
4.5 It verifies and explains that a moderate WSD annealing ratio could get the lowest loss.
In the case of WSD learning rate schedule, it is crucial to ascertain the optimal annealing ratio for training steps. Prior research by Hägele et al. (2024) has established the existence of an optimal annealing ratio within the WSD scheduler. As illustrated in experimental results of their study, excessively high or low annealing ratios lead to sub-optimal model performance. This phenomenon can be further elucidated through our proposed equation.
Theoretically, a high annealing ratio results in a significant reduction of the forward training area $S_{1}$, while only marginally increasing the annealing area $S_{2}$. Conversely, an excessively low annealing ratio leads to under-annealing, characterized by a diminutive $S_{2}$ and consequently get a high final loss. Our scaling law function builds a trade-off relationship between the forward area $S_{1}$ and annealing area $S_{2}$ about the annealing ratio.
As depicted in Fig. 11, we utilize the fitted scaling law function to draw the final loss across various annealing ratios and total training steps. The prediction results present parabola-like curves, and align well with the actual experimental outcomes by previous works. It suggests that a moderate WSD annealing ratio is optimal which could make a moderate $S_{1}$, and then maximize $S_{1}$ and $S_{2}$ from a global view, thereby minimizing the overall validation loss. Moreover, our equation directly help predict an optimal annealing ratio for different total steps without experiments, which saves a lot of resources.
@@
4.6 It verifies and explains that the optimal annealing function in WSD LRS depends on the annealing ratio.
@@
Figure 12: The predicted (left) and true loss (right) of cosine and 1-sqrt annealing method at different annealing ratios. Experimental results (right), aligned with our prediction (left), refute the previous finding “the order and results of different annealing hold across settings” (Hägele et al., 2024).
@@
In the context of the WSD LRS, the selection of the annealing method in the annealing stage is also pivotal to optimize the training process. Hägele et al. (2024) conclude that the 1-sqrt annealing (refer to Appendix E for 1-sqrt function and curve) yields a lower final loss compared to the other annealing methods (e.g. cosine). They claim that the conclusion holds true across different annealing ratios.
@@
However, as we predict using our equation (Fig. 12(a)), it indicates that 1-sqrt annealing does get a lower loss than cosine annealing in 10% and 20% annealing ratios, but performs much worse than cosine annealing in 50% annealing ratio.
@@
To verify whether the predictions from our equation are accurate, we conduct experiments by training models using different annealing methods and ratios within a fixed 50K total steps. As illustrated in Fig. 12(b), at a 10% or 20% annealing ratio, the 1-sqrt method outperforms the cosine method, whereas at a 50% annealing ratio, the latter method exhibits a lower final loss. The true experimental results align quite well with our prediction, which also overturns some of the conclusions made by previous works. We conclude that the optimal annealing function in WSD LRS depends on the annealing ratio.
@@
Our scaling law function provides an explanatory framework for these observations. We draw the LR curves of 1-sqrt and cosine annealing in Appendix E. At 10% annealing ratio, although the forward area $S_{1}$ of the cosine method is slightly larger than that of the 1-sqrt method, the larger annealing area $S_{2}$ of the 1-sqrt method plays a more critical role in reducing the overall final loss. However, as the annealing ratio increases, the difference of $S_{1}$ between two LRS gradually becomes larger and larger, till breaking the delicate balance between $S_{1}$ and $S_{2}$ at 50% annealing ratio, resulting in a lower final loss for the cosine method. This relationship underscores the importance of carefully selecting the annealing strategy to optimize model training outcomes within the WSD scheduler. Still, our equation can help predict a better annealing method without experiments, which saves a lot of resources.
@@
4.7 It verifies and explains that in continual pre-training, the higher max learning rate to re-warmup, the higher the initial peak loss will be, and then the more sharply it will decrease.
@@
In continual pre-training (CPT), the learning rate scheduler is usually set as re-warmup to a new max LR at the beginning. By many experiments, Gupta et al. (2023) concludes that the higher max learning rate to re-warmup, the higher the initial peak loss will be, and then the more sharply it will decrease.
According to our scaling law function  [^4] Strictly speaking, continual pre-training process often include LR re-warmup as well as data distribution shift. Here we primarily research on the condition where there is no distribution shift between two training stages. The conclusions transfer across most cases because the loss change brought by LR re-warmup is significantly larger than the loss change brought by data distribution shift (Gupta et al., 2023; Ibrahim et al., 2024)., in the re-warmup process, the annealing area $S_{2}$ will reduce to a negative value ($S_{2}<0$) and thus the validation loss increases. The higher max LR in re-warmup, the annealing area $S_{2}$ becomes more negative and thus there would be a higher peak loss. But still, higher max LR could make the forward area $S_{1}$ grow faster and the loss decreases more sharply after re-warmup. We use the fitted equation to predict the continual pre-training process with different max LR as shown in Fig. 13(a). The predicted loss curves reproduce a quite similar phenomenon with previous works (Gupta et al., 2023).
There is a more profound strategy using our equation in CPT. As shown in Fig. 13(a), after ensuring total steps during CPT, we can apply our equation to predict a better max LR and scheduler to get the lowest final loss without experiments, which saves a lot of resources.
Meanwhile, how many steps to re-warmup is another important issue in the continual pre-training.  Gupta et al. (2023) find that the longer re-warmup steps could smooth the transition of loss curve but the number of re-warmup steps does not significantly influence the final validation loss. We use the fitted equation to predicted the continual pre-training dynamics with different re-warmup steps. The results, shown in Fig. 13(b), present a good alignment with previous works (Gupta et al., 2023).
Based on our theory, given the fixed max LR, when the re-warmup steps are longer, the annealing area decreases more slowly and the loss curve rises more smoothly, but both final $S_{1}$ and $S_{2}$ are quite stable across different re-warmup steps. First, the annealing area $S_{2}$ of different re-warmup steps are very close due to the same max LR and the same min LR. Besides, though different re-warmup steps bring in temporary distinct losses, re-warmup only cover a small percentage compared with all training steps. Thus, the forward area $S_{1}$ is also close across different re-warmup steps, resulting in the close overall loss across different steps of re-warmup.
@@ 5.1 Reduction to Chinchilla Scaling Law
Our scaling law function can predict the full loss curve across any given learning rate scheduler. In this section, we show that our equation has no contradiction with typical scaling law, and it is a generalized form of the chinchilla scaling law (Hoffmann et al., 2022). That is to say, all the final loss points of different total training steps following our equation should also follow a power-law relationship. We prove this by dividing into two conditions: (1) constant LRS, and (2) other LRS.
Constant LRS.
In the case of constant learning rate scheduler, the annealing area $S_{2}$ is always zero and the forward area $S_{1}=\eta_{max}\cdot s$, where $s$ is the step, and $\eta_{max}$ is the maximal learning rate which is constant over steps. Thus, the whole train loss curves becomes:
@@
which can be easily reduced to the chinchilla scaling law equation.
Other LRS.
For other learning rate schedulers, we adopt a method based on statistics to show that our scaling law function can be reduced to the chinchilla scaling law. Specifically, we check whether chinchilla scaling law fits well the endpoints of loss curves predicted and generated by our scaling law. The parameter tuple of our equation is $(L_{0},A,C,\alpha)$. We then randomly sample 1000 sets of parameter tuples in some uniform distributions: $L_{0}\sim U(1,3)$, $A\sim U(0.3,0.5)$, $C\sim U(0.2,0.6)$, $\alpha\sim U(-0.6,-0.4)$. Each parameter tuple could be seen as the fitting result of a distinct set of experimental setups  [^5] It’s worth noting that some of these sampled parameter tuples might not be reasonable or likely to happen in real-world scenarios, but we choose to keep them nonetheless.. (e.g. dataset, batch size, model size, etc.). For each generated parameter tuple, we apply our equation to predict the final loss of different total training steps range from 5K steps to 60K steps. We conduct the prediction on two LRS including cosine and WSD (10% annealing ratio). The predicted final loss points are used to fit the chinchilla equation through minimizing the Huber loss. Fitting examples are shown in Fig. 14.
@@
Figure 14: Chinchilla scaling law fits well the validation loss endpoints predicted by our formulation, taking cosine LRS (on the left) and WSD LRS (on the right) as examples.
LR Scheduler $\text{mean}(R^{2})$uparrow $\text{std}(R^{2})$downarrow Huber loss $\downarrow$
Cosine 0.972 0.056 0.00017
WSD 0.979 0.053 0.00013
Table 1: Mean and standard deviation of $R^{2}$, as well as the mean Huber loss of $1000$ randomly generated parameter tuples.
We calculate the mean and standard deviation of $R^{2}$ and Huber loss among 1000 random parameter tuples. As shown in Table 1, the random selected parameter tuples have low Huber loss and high $R^{2}$ value. The experimental results demonstrate that chinchilla scaling law fits well on the data predicted and generated by our scaling law function. That is, our scaling law with LR annealing can be reduced to chinchilla scaling law.
5.2 Scaling Law Fitting Democratization
We extremely democratize the fitting and prediction of scaling law by greatly enlarging its scope of applicability. Adopting our scaling law with LR annealing, we can collect thousands of fitting points in one full loss curve during training, while typical scaling law can only collect one endpoint from a full loss curve.
For a more direct comparison with chinchilla scaling law (Hoffmann et al., 2022), we suppose a scenario where it requires 100 fitting points to get the parameters of scaling laws. We assume the distance between each point as $K$. We compute the required training steps using different approaches as follows:
  • Adopting chinchilla scaling law, typical cosine LRS requires total steps of at least $1K+2K+3K+\cdots+100K=5050K$.
@@ 5.2 Scaling Law Fitting Democratization
  • Adopting our scaling law, all we need is only one training curve with moderate total steps (and the number of fitting points is far more than $100$), such as one curve with $50K$ steps  [^6] The empirical rule that more fitting points always achieve better fitting results always holds true. Our equation can also use more points and LRS for fitting, such as $30K$ constant + $70K$ cosine. Nevertheless, we can collect far more fitting points than the typical scaling law with significantly fewer training steps..
We present a comparison of the computational costs associated with different laws and LRS in Table 2. Our findings indicate that our proposed equation enables us to fit loss with less than 1% of the computational cost required by the chinchilla scaling law. Furthermore, our scaling law with LR annealing, can be universally applicable to other LRS by fitting any given LRS (except constant), thereby conserving additional computational resources. This approach extremely democratizes the process of fitting and predicting scaling laws in LLM pre-training, paving the way for a more environmentally friendly and carbon-efficient methodology.
@@
Table 2: The comparison of computational cost for fitting, assuming $100$ fitting points as an example.
@@ 6.1 The impact of Decay Factor $\lambda$
The decay factor $\lambda$ in our equation plays a crucial role to indicate the information retaining degree in LR annealing. We set $\lambda$ as $0.999$ in our all experiments. We explore the difference from another decay factor $\lambda=0.99$. We fit and get different equations for different $\lambda$. We compare them (1) on the predicted loss curves for 1-square and 1-sqrt annealing methods, and (2) on the predicted loss curves in different annealing ratios of WSD LRS (cosine annealing).
The results, illustrated in Fig. 15 and 16, reveal several key insights into the impact of decay factor:
@@
More essentially, the selection of $\lambda$ decides the balance point of $S_{1}$ and $S_{2}$. For example, $\lambda=0.999$ means that, LR annealing only retain the information of previous approximately $\frac{1}{1-\lambda}=1000$ steps, which can be seen as the window size of LR annealing momentum. The window size could be very close to the optimal annealing steps. After reaching window size, $S_{2}$ increases very slowly, with the cost of large decrease of $S_{1}$.
The analyses above underscore the sensitivity of our scaling law function to the choice of decay factor and highlights the importance of selecting a decay factor that aligns closely with empirical data to ensure the accuracy of predictions. We recommend that the future developers try different $\lambda$ for their own setups  [^7] Actually, $\lambda$ can be fitted as a parameter, instead of a hyper-parameter requiring manual tuning. We regard $\lambda$ as a hyper-parameter because $\lambda=0.999$ performs well in our most experiments. Besides, fitting with $\lambda$ could bring in additional time complexity due to the recomputation of $S_{2}$..

### v2

@@ 4 Takeaways: Experimental Findings Verification and Explanation
Recent studies have shown that WSD LRS (Hu et al., 2024) and multi-step cosine LRS (DeepSeek-AI, 2024) result in lower loss compared to the traditional cosine LRS. Our experiments also support this finding (refer to the ground-truth loss in Fig. 3(b)3(c)3(d)). We validate and elucidate this finding using our proposed equation. Fig. 11 shows the learning rate curve (left) and the predicted loss drop (right) for different LRS. The results suggest that for WSD and multi-step cosine LRS, the negative $S_{2}$-item ($-C\cdot S_{2}$) is slightly larger than that of the cosine LRS, whereas the $S_{1}$-item ($A\cdot S_{1}^{-\alpha}$) is significantly lower. Specifically, both the WSD LRS and multi-step cosine LRS unintentionally employ strategies that marginally reduces $S_{2}$ but substantially increases $S_{1}$, leading to an overall decrease in validation loss.
4.5 Determining Optimal Annealing Ratio of WSD LRS.
In the case of WSD LRS, it is crucial to ascertain the optimal annealing ratio for training steps.  Hägele et al. (2024) found that there is an optimal annealing ratio for WSD LRS, i.e. both excessively high or low annealing ratios lead to sub-optimal model performance. This phenomenon can be further elucidated through our proposed equation. Specifically, a high annealing ratio results in a significant reduction of the forward area $S_{1}$ while a low annealing ratio leads a diminished annealing area $S_{2}$. Our scaling law equation describes the trade-off between the forward area $S_{1}$ and the annealing area $S_{2}$ about the annealing ratio.
Fig. 12 depicts the final loss predicted by our equation for various annealing ratios and total training steps. The predictions form parabola-like curves, and align well with the actual experimental results reported in previous studies. This suggests that a moderate annealing ratio, typically around 10% to 20%, is optimal, as it balances $S_{1}$ and $S_{2}$ to maximize their combined effect, thereby minimizing the overall validation loss. Moreover, our equation can directly predict the optimal annealing ratio for different total steps without large-scale experiments, which saves a lot of resources.
@@
4.6 Determining the optimal annealing function in WSD LRS.
@@
Figure 13: The predicted (left) and true loss (right) of cosine and 1-sqrt annealing method at different annealing ratios. Experimental results (right), aligned with our prediction (left), refute the claim “the order and results of different annealing hold across settings” (Hägele et al., 2024).
@@
For the WSD LRS, the selection of the annealing function is another important factor for obtaining further lower loss values. Hägele et al. (2024) conclude that the 1-sqrt annealing (see Appendix 10 for the detailed formula) yields lower final losses compared to the other annealing methods (e.g. cosine). They claim that this conclusion is true across different annealing ratios.
@@
However, our equation suggests different results (see Fig. 13(a)). The 1-sqrt annealing approach does get a lower loss than the cosine annealing approach when using small annealing ratios (e.g. 10%), but it performs much worse than the cosine annealing approach when using 50% annealing ratio.
@@
To verify our prediction, we conduct experiments by training models using different annealing methods and ratios within 50K total steps. As illustrated in Fig. 13(b), at a 10% annealing ratio, the 1-sqrt annealing method obtains lower final losses compared to the cosine annealing method, whereas at a 50% annealing ratio, the cosine annealing method obtains a lower final loss compared to the 1-sqrt annealing method. The experimental results align quite well with our prediction, which also overturns some of the claims made by previous works. We conclude that the optimal annealing function in WSD LRS depends on the annealing ratio.
@@
Our scaling law equation provides an explanatory framework for these empirical observations. We draw the LR curves associated with the 1-sqrt and cosine annealing approaches in Appendix 10. When using an small annealing ratio, the forward area $S_{1}$ of the cosine annealing method is slightly larger than that of the 1-sqrt annealing method, while the annealing area $S_{2}$ of the 1-sqrt method plays a more dominate role in decreasing the final loss. As the annealing ratio increases, the difference of $S_{1}$ between two annealing methods becomes larger and larger. Therefore the forward area $S_{1}$ gradually takes the dominate role. The delicate balance between $S_{1}$ and $S_{2}$ breaks at 50% annealing ratio, resulting in a lower final loss for the cosine annealing function.
The above analysis underscores the importance of carefully selecting the annealing strategy for the WSD LRS to optimize model training outcomes. Our equation can help predict a better annealing function without experiments, which saves a lot of resources.
@@
4.7 It verifies and explains that in continual pre-training, a higher max learning rate during re-warmup leads to a higher initial peak in the loss, followed by a more rapid decrease.
In continual pre-training (CPT) settings, the LRS typically involves re-warmup to a new max LR at the start of training. Through numerous experiments, Gupta et al. (2023) concludes that a higher max LR during re-warmup results in a higher initial peak loss, followed by a more rapid decrease.
According to our scaling law formulation  [^3] Strictly speaking, continual pre-training process often include LR re-warmup as well as data distribution shift. Here we primarily research on the setting that there is no distribution shift between two training stages. The conclusions transfer across most cases because the loss change brought by LR re-warmup is significantly larger than the loss change brought by data distribution shift (Gupta et al., 2023; Ibrahim et al., 2024)., in the LR re-warmup process, the annealing area $S_{2}$ will reduce to a negative value ($S_{2}<0$) and thus the validation loss increases. A higher max LR in re-warmup lead to a lower annealing area $S_{2}$, and thus there would be a higher peak loss. Moreover, a higher max LR also leads to a faster growth of the forward area $S_{1}$ after re-warmup. We use the fitted equation to predict the continual pre-training process with different max LR as shown in Fig. 14(a). The predicted loss curves reproduce a quite similar phenomenon with previous works (Gupta et al., 2023).
There is a more profound strategy using our equation in CPT. As shown in Fig. 14(a), after the total steps of CPT is determined, we can apply our equation to predict a better max LR and scheduler to get the lowest final loss without experiments, which saves a lot of resources.
Meanwhile, how many steps to re-warmup is another important issue in the continual pre-training.  Gupta et al. (2023) find that longer re-warmup steps smooth the transition of loss curve but the number of re-warmup steps does not significantly influence the final validation loss. We use the fitted equation to predict the continual pre-training dynamics with different re-warmup steps. The results, shown in Fig. 14(b), present a good alignment with previous observations (Gupta et al., 2023).
Based on our theory, given a fixed max LR, longer re-warmup steps lead to a slower decreases of the annealing area, resulting in a smoother raise in the loss curve. However, both the final values of the forward area ($S_{1}$) and the annealing area ($S_{2}$) remain relatively stable across different re-warmup steps. The annealing area ($S_{2}$) corresponding to different re-warmup steps are very close since the max and min LR remain the same. Besides, although different re-warmup steps lead to temporary distinct loss fluctuations, re-warmup only accounts for a small portion of the whole training process. Thus, the forward area $S_{1}$ is also close across different re-warmup steps, resulting in the close overall loss across different steps of re-warmup.
@@ 5.1 Reduction to Chinchilla Scaling Law
Our scaling law equation can predict the full loss curve across any given LRS. In this section, we show that our equation has no contradiction with traditional scaling laws, and it is a generalized form of the Chinchilla scaling law (Hoffmann et al., 2022). Specifically, all the final loss values for different total training steps following our equation should also follow a power-law relationship. We prove this by dividing two conditions: (1) constant LRS, and (2) other LRS.
Constant LRS. In the case of a constant LRS, the annealing area $S_{2}$ is always zero and the forward area $S_{1}=\eta_{max}\cdot s$, where $s$ is the step, and $\eta_{max}$ is the constant maximal LR. Thus, the whole loss curve formula becomes:
@@
which aligns with the Chinchilla scaling law equation.
Other LRS. For non-constant LRS, we use a statistical approach to show that our equation can be reduced to the Chinchilla scaling law. Specifically, we verify whether the Chinchilla scaling law adequately fits the endpoints of loss curves predicted by our equation. The parameter tuple of our equation is $(L_{0},A,C,\alpha)$. We then randomly sample 1000 sets of parameter tuples from some uniform distributions: $L_{0}\sim U(1,3)$, $A\sim U(0.3,0.5)$, $C\sim U(0.2,0.6)$, $\alpha\sim U(-0.6,-0.4)$. Each parameter tuple could be seen as the fitting result of a distinct set of experimental setups  [^4] It’s worth noting that some of these sampled parameter tuples might not be reasonable or likely to be observed in real experiments, but we choose to keep them nonetheless.. (e.g. dataset, batch size, model size, etc.). For each generated parameter tuple, we apply our equation to predict the final loss of different total training steps on two LRS including cosine and WSD (10% annealing ratio), ranging from 5K to 60K steps. The predicted loss values on final steps are used to fit the chinchilla equation. The fitting results are shown in Fig. 15.
Table 1: Mean and standard deviation of $R^{2}$, as well as the mean Huber loss of 1,000 randomly generated parameter tuples.
LR Scheduler mean($R^{2}$) $\uparrow$ std($R^{2}$) $\downarrow$ Huber Loss $\downarrow$
Cosine 0.972 0.056 0.00017
WSD 0.979 0.053 0.00013
@@
Figure 15: The Chinchilla scaling law equation fits well on the final validation losses predicted by our formulation, taking cosine LRS (on the left) and WSD LRS (on the right) as examples.
Each parameter tuple represents a synthetic fitting result corresponding to a distinct set of experimental setups (e.g. dataset, model size, etc.). For each sampled parameter tuple, we apply our equation to predict the final loss for different total training steps with both cosine and WSD LRS, and then employ the predicted losses to fit the Chinchilla scaling law. We calculate the mean and standard deviation of $R^{2}$ values for each fit. The results in Table 1 demonstrate that Chinchilla scaling law fits well on the data predicted by our scaling law equation. Thus, our equation can be considered a generalization that can be reduced to the Chinchilla scaling law.
5.2 Scaling Law Fitting and Prediction Democratization
Our scaling law equation allows us to utilize all loss values from a full loss curve during training, while traditional scaling laws can only collect a single data point from the full loss curve. This feature allows us to fit scaling laws with much less cost. For a direct comparison, we compare the computational efficiency of our approach and the Chinchilla scaling law (Hoffmann et al., 2022). Specifically, we assume and evaluate the computational cost of obtaining 100 fitting points for each scaling law equation with an step interval of $K$:
  • Adopting Chinchilla scaling law, typical cosine LRS requires total steps of at least $1K+2K+3K+\cdots+100K=5050K$;
@@ 5.2 Scaling Law Fitting and Prediction Democratization
  • Adopting our scaling law, all we need is only one or two training curves with moderate total steps, such as one curve with $50K$ steps under cosine LRS. Further, the number of loss values we can get for fitting our equation is far more than $100$. Note that more fitting points generally achieve more accurate fitting results. Our equation can also utilize loss curves from various different LRS (e.g. $20K$ constant + $30K$ cosine)
Table 2: The comparison of computational cost for fitting different scaling law equations.
@@
We present a comparison of the computational costs associated with different laws and LRS in Table 2. The results indicate that our proposed equation uses less than 1% of the computational cost required by the Chinchilla scaling law. Further, our scaling law with LR annealing, can be universally applied to predict full loss curves for unseen LRS, thus conserving even more computational resources. This approach significantly democratizes the study of scaling laws in LLM pre-training, paving the way for a more environmentally friendly and carbon-efficient methodology.
@@ 6.1 The impact of Decay Factor $\lambda$
The decay factor $\lambda$ in our equation indicates the information retaining degree in LR annealing. We set $\lambda$ as $0.999$ in our all experiments. We explore the difference from another decay factor $\lambda=0.99$. We fit and get different equations for different $\lambda$. We compare them (1) on the predicted loss curves for 1-square and 1-sqrt annealing methods, and (2) on the predicted loss curves in different annealing ratios of WSD LRS (cosine annealing).
The results, illustrated in Fig. 16 and 17, reveal several key insights into the impact of decay factor:
@@
More essentially, the selection of $\lambda$ decides the balance point of $S_{1}$ and $S_{2}$. For example, $\lambda=0.999$ means that LR annealing only retain the information of previous approximately $\frac{1}{1-\lambda}=1000$ steps, which can be seen as the window size of LR annealing momentum. The window size could be very close to the optimal annealing steps if LR changes smoothly over steps. After reaching window size, $S_{2}$ increases very slowly, with the cost of large decrease of $S_{1}$.
The analyses above highlights the importance of selecting a decay factor that aligns closely with empirical data to ensure the accuracy of predictions. We recommend that the future developers try different $\lambda$ for their own setups  [^5] Actually, $\lambda$ can be fitted as a parameter, instead of a hyper-parameter requiring manual tuning. We regard $\lambda$ as a hyper-parameter because $\lambda=0.999$ performs well in our all experiments. Besides, fitting with $\lambda$ could bring in additional time complexity due to the recomputation of $S_{2}$..